# ProxyLM: Predicting Language Model Performance on Multilingual Tasks via Proxy Models

### v1

@@ A.3 Hyper-parameters
Each fine-tuning and evaluation for LMs is done with an NVIDIA Tesla V100 32GB GPU. The hyper-parameters used during fine-tuning from the MT560 Gowda et al. (2021) and NusaTranslation Cahyawijaya et al. (2023) datasets are listed in Table 10, 11, 12, and 13 for SMaLL100, M2M100, NLLB, and Transformer models, respectively.
@@
Each regressor is trained on an AMD Ryzen Threadripper 2990WX with 128 GB of RAM and 16 threads. Regressors’ hyper-parameters used are provided in Table 14, 15, 16, and 17 for XGB, Poly2/Poly3, LGBM, and MF, respectively. These hyper-parameters were obtained based on the best cross-validation RMSE score using 10 folds.
@@
Appendix B More Detailed Results
We provide detailed visualizations of the results of XGBoost with ProxyLM Ensemble based on multiple language groupings in Figure 5 - 10 for English-centric result, and Figure 11 - 14 for Many-to-Many Languages result. Each language groupings plot comprises multiple subplots, including (a) vitality class, (b) Joshi class, (c) language family, and (d) individual languages. The mapping of vitality, Joshi class, and language family follows the classifications in Table 8 and 9.
| Hyper-parameter | MT560 | NusaTranslation |
@@ Appendix B More Detailed Results
| Hyper-parameter | MT560 | NusaTranslation |
@@
Appendix C Further Analysis
C.1 Model Feature Importance
We provide feature importance scores of XGBoost with ProxyLM Ensemble for the random MT560 experiment in Figure 15 and 16. Each combination consists of one most influential feature followed by others with marginal contributions to the model, each with an importance score of 0.12 or less. We observe that proxy models are always the most influential features in prediction.
| Hyper-parameter | MT560 | NusaTranslation |
@@ C.1 Model Feature Importance
| Hyper-parameter | MT560 | NusaTranslation |
@@
| Hyper-parameter | NusaTranslation | MT560 | ||

### v2

@@ A.3 Hyper-parameters
Each regressor is trained on an AMD Ryzen Threadripper 2990WX with 128 GB of RAM and 16 threads. Regressors’ hyper-parameters used are provided in Table 14, 15, 16, and 17 for XGBoost, Poly2/Poly3, LGBM, and MF, respectively. These hyper-parameters were obtained based on the best cross-validation RMSE score using 10 folds.
@@
| Hyper-parameter | English-centric | Many-to-Many Langs. |
@@
| Hyper-parameter | English-centric | Many-to-Many Langs. |
@@
A.4 Regressor Dataset Sizes
We provide the details of the regressor’s training and test set size in Table 18.
Appendix B More Detailed Results
We provide detailed visualizations of the results of XGBoost with ProxyLM Ensemble based on multiple language groupings in Figures 5-10 for English-centric result, and Figures 11-14 for Many-to-Many Languages result. Each language groupings plot comprises multiple subplots, including (a) vitality class, (b) Joshi class, (c) language family, and (d) individual languages. The mapping of vitality, Joshi class, and language family follows the classifications in Table 8 and 9.
Appendix C Feature Importances
We provide feature importance scores of XGBoost with ProxyLM Ensemble for the random English-centric experiment in Figure 15 and 16. Each combination consists of one most influential feature followed by others with marginal contributions to the model, each with an importance score of 0.12 or less. We observe that proxy models are always the most influential features in prediction.
Appendix D License for Artifacts
We discuss the license or terms for the use of any artifacts we use in Table 19.
| Hyper-parameter | English-centric | Many-to-Many Langs. |
@@ Appendix D License for Artifacts
| Hyper-parameter | English-centric | Many-to-Many Langs. |
@@
| Hyper-parameter | Many-to-Many Langs. | English-centric | ||
@@
| Experimental Settings | Train Size | Test Size |
| Random (English-centric) | 829 | 356 |
| Random (Many-to-Many Langs.) | 156 | 68 |
| Unseen | 1,084 | 101 |
| Cross-Dataset | 1,185 | 224 |
| Datasets | URL Link | License |
| MT560 Gowda et al. (2021) | https://opus.nlpl.eu/MT560 | Unknown |
| FLoRes Costa-jussà et al. (2022b) | Muennighoff/flores200 | CC-BY-SA 4.0 |
| NusaTranslation Cahyawijaya et al. (2023) | https://huggingface.co/datasets/indonlp/nusatranslation_mt | Apache 2.0 |